OpenStudy (anonymous):

Write an equation of the line in point slope form that passes through the two given points. (-2,15) and (9, -18)?

OpenStudy (dobby1):

\[m=\frac{ -18-15 }{ 9-(-2) }\]

OpenStudy (dobby1):

\[m=\frac{ -33 }{ 11 }\]

OpenStudy (dobby1):

point sloap form looks like \[y-y _{1}=m(x-x _{1})\]

OpenStudy (dobby1):

\[y-15=\frac{ -33 }{ 11 }(x-(-2)\] -33/11=-3

OpenStudy (dobby1):

\[y-15=-3(x-(-2))\]

OpenStudy (dobby1):

@canadalover_ @mrivera3113

OpenStudy (dobby1):

Could i have a medal?

OpenStudy (dobby1):

THANKS

OpenStudy (dobby1):

TANKS

OpenStudy (dobby1):

oops THANKS